# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ Appendix A Missing Details in Preliminaries
Example A.1 (Query Classes).
@@
Definition A.2 (Tree Decomposition).
@@
Definition A.3 (The restriction $Q_{\bm{Y}}$ of a query $Q$).
@@
Definition A.4 (Fractional Hypertree Width).
@@
Conjecture A.6 (OuMv (Henzinger et al., 2015)).
@@
Appendix B Missing Details in Section 5
In this section, we explain in detail our results on IVM in the insert-only setting, where no deletes are allowed. The main goal is to prove Theorem 3.1 (as well as Theorem 3.3). The results in this section also serve as building blocks for our results on the insert-delete setting in Section D.
B.1. Preliminaries: The Query Decomposition Lemma
We prove the query decomposition lemma (Lemma 17) and illustrate it by an example. This lemma is the core of the proof of Lemma 5.1, which in turn is a key ingredient in the proof of Theorem 3.1. Recall notation from Section A.
@@ B.1. Preliminaries: The Query Decomposition Lemma
Lemma B.1 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
@@
(23) $\sum_{y\in\text{\sf Dom}(Y)}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|\sigma_{Y=y}R(\bm{X})|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
@@
(33) $\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}\left(\sum_{y\in\text{\sf Dom}(Y)}|\sigma_{Y=y}R(\bm{X})|\right)^{\lambda_{R(\bm{X})}}$ $=$
(38) $\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
@@
B.2. Proof of Lemma 5.1
Theorem 3.1 gives an upper bound on the amortized update time for IVM in the insert-only setting in terms of the fractional hypertree width of the query $Q$. Before we show how to achieve this bound in Sections B.3 and B.4, we prove in this section a weaker bound that is based on the AGM-bound of $Q$. In particular, we prove Lemma 5.1, which says the following. Given a query $Q$ (whose input relations are initially empty) and a stream of $N$ single-tuple inserts into $Q$, we can update $Q$ after every insert and report the new output tuples $\delta Q$ in a total time for all updates/reports together that does not exceed the AGM bound of $Q$. Note that here we are referring to the final AGM bound of $Q$ after all inserts have been performed.
@@
$\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(\tau)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$$\leq$
$\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$$\leq$
$\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}|^{\lambda_{R(\bm{X})}}$ $=\theta$
@@
B.3. Upper bound for $\text{IVM}^{+}[Q]$
@@
[…]
@@
The case of $\text{IVM}^{+}[Q]$: Let $(\mathcal{T},\chi)$ be a tree decomposition of $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$. (Recall notation from Section A.) Fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. Given a node $t\in V(\mathcal{T})$, we will use $\text{\sf parent}(t),\text{\sf children}(t),\text{\sf ancestors}(t)$ and $\text{\sf decendants}(t)$ to refer to the parent, children, ancestors and descendants of $t$ in $\mathcal{T}$, respectively, under the fixed orientation of $\mathcal{T}$. For every node $t\in V(\mathcal{T})$, let $\bm{X}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\chi(t)$ be the variables of $t$ and $Q_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{\bm{X}_{t}}$ be the restriction of $Q$ to $\bm{X}_{t}$, given by Definition A.3. We call $Q_{t}$ a bag query. Let $\bm{X}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf vars}(Q)$. By definition, the query $Q$ is equivalent to the join of the outputs of all bag queries:
@@
B.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Theorem 3.1 for the $\text{IVM}^{+}_{\delta}[Q]$ problem. In particular, we explain how to modify the algorithm and proof from Section B.3, which was originally meant for the $\text{IVM}^{+}[Q]$ problem. Just like we did in Section B.3, here also we will construct an optimal tree decomposition of $Q$ that minimizes the fractional hypertree width, and we will maintain materialized relations for every bag of the tree decomposition. However, suppose that we have an insert of a tuple $\bm{t}$ into a relation $R$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this particular insert. We can no longer start our constant-delay enumeration from the root of the tree decomposition, as we did in Section B.3. Instead, we have to find the bag $B$ that contains $R$, select $\bm{t}$, and start our enumeration from $B$, as if $B$ were the root of the tree decomposition. In order to be able to enumerate the output of $Q$ starting from any bag $B$ of the tree decomposition, all bags need to be fully calibrated with each other. In particular, in Section B.3, we only calibrated bags in a bottom-up fashion where each bag was calibrated with its children bags. But here, we have to do full calibration consisting of two phases: One bottom-up phase where we calibrate each bag with its children, followed by a top-down phase where we calibrate each bag with its parent. These two calibrations are similar to the two phases of the Yannakakis algorithm (Yannakakis, 1981).
@@ B.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
[…]
@@
The case of $\text{IVM}^{+}_{\delta}[Q]$: We will use the same notation as in the proof of Theorem 3.1 for $\text{IVM}^{+}[Q]$ in Section B.3. In particular, here also we will fix an optimal tree decomposition $(\mathcal{T},\chi)$ of $Q$ and we will fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. In addition to $Q^{\prime}_{t}$ and $P^{\prime}_{t}$ given by (41) and (42), we will maintain two additional materialized relations $Q_{t}^{\prime\prime}$ and $P_{t}^{\prime\prime}$ for every node $t\in V(\mathcal{T})$. However, this time, we will define them inductively starting from the root $r$ all the way to the leaves of $\mathcal{T}$: (Recall that for the root $r$, we have $\bm{Y}_{r}=\emptyset$.)
@@
Now suppose we have an insert of a tuple $\bm{t}$ into a relation $R(\bm{Z})$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this insert. We find the node $t$ in $\mathcal{T}$ where $\bm{Z}\subseteq\chi(t)$. The existence of such a node $t$ is guaranteed by the definition of a tree decomposition (Definition A.2). Now, we re-orient the tree $\mathcal{T}$ so that $t$ becomes the root. We enumerate output tuples $\bm{x}_{t}$ of $Q^{\prime\prime}_{t}$ that join with $\bm{t}$. For each such tuple $\bm{x}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime\prime}_{c}$ for every child $c$ of $t$. (These are the children of $t$ under the new orientation of the tree $\mathcal{T}$). By (44), we guarantee that for each tuple $\bm{x}_{t}$ and each child $c$, there must be at least one tuple $x_{c}$ in $Q^{\prime\prime}_{c}$ that joins with $\bm{x}_{t}$. We continue this process inductively all the way to the leaves of $\mathcal{T}$. ∎
Example B.3 (for Theorem 3.1).
@@
B.5. Lower bound for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$
@@
[…]
@@
Appendix C Missing Details in Section 6
In this section, we review some necessary background on intersection joins in the static setting. Let $Q$ be a query, $\overline{Q}$ be its time extension, and $\widehat{Q}$ be its multivariate extension. Suppose we want to evaluate $\overline{Q}$ over a given database instance $\overline{\mathcal{D}}$. This evaluation problem can be reduced (Khamis et al., 2022) to the evaluation of $\widehat{Q}$ over some database instance, $\widehat{\mathcal{D}}$, that satisfies $|\widehat{\mathcal{D}}|=\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$ and can be constructed in time $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$. Moreover, this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown (Khamis et al., 2022) by a backward reduction from each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail below. Recall that both reductions are in the context of static evaluation. However later on in Section D, we will rely on both reductions in order to prove upper and lower bounds on the IVM problem in the insert-delete setting.
C.1. A reduction from $\overline{Q}$ to $\widehat{Q}$
@@
Example C.2 ().
Suppose we want to solve the query $\overline{Q}_{\triangle}$ from Example 6.3. We can start by constructing a segment tree for the interval variable $[Z]$ and replacing each value $[z]$ of $[Z]$ by the components in its canonical partition $\text{\sf CP}_{N}([z])$. We can then use Lemma C.1 to look for intersections. In particular, let’s first try to find triples of segment-tree nodes $(z^{R},z^{S},z^{T})$ from $R,S,T$ respectively that appear on the same root-to-leaf path in the specific order $(z^{R},z^{S},z^{T})$. This can only happen if the bitstring $z^{R}$ is a prefix of $z^{S}$ which in turn is a prefix of $z^{T}$. To test this prefix pattern, we can break each $z^{S}$ into a concatenation two strings $z^{S}=z^{S}_{1}\circ z^{S}_{1}$ in all possible ways and similarly break $z^{T}$ into a concatenation of three string $z^{T}=z^{T}_{1}\circ z^{T}_{2}\circ z^{T}_{3}$ and then check whether $z^{R}=z^{S}_{1}=z^{T}_{1}$ and $z^{S}_{2}=z^{T}_{2}$. In particular, we construct the following input relations:
@@
$\widehat{T}_{123}$ =$=\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge$
$\quad\quad\quad\quad(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
@@
(47) $\widehat{Q}_{123}(Z_{1},Z_{2},Z_{3},A,B,C)=\\ \widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
@@
Definition C.3 (Canonical partition of a relation $\text{\sf CP}_{N}^{(i)}(R)$).
@@
(49) $\text{\sf CP}_{N}^{(i)}([z],x_{1},\ldots,x_{m})\stackrel{{\scriptstyle\text{def}}}{{=}}\\ \{(z_{1},\ldots,z_{i},x_{1},\ldots,x_{m})\mid(z_{1}\circ\cdots\circ z_{i})\in\text{\sf CP}_{N}([z])\}$
@@
Definition C.4 (Canonical partition of a database instance $\text{\sf CP}_{N}(\overline{\mathcal{D}})$).
@@
Theorem C.5 (Implicit in (Khamis et al., 2022)).
@@
[…]
@@
C.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
In the previous Section C.1, we saw how to reduce the static evaluation of $\overline{Q}$ to $\widehat{Q}$. We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Khamis et al., 2022). Later on, we will use this backward reduction to prove lower bounds on the IVM problem in Section D.4.
Example C.6 ().

### v3

@@ Appendix B Missing Details in Preliminaries
Example B.1 (Query Classes).
@@
Definition B.2 (Tree Decomposition).
@@
Definition B.3 (The restriction $Q_{\bm{Y}}$ of a query $Q$).
@@
Definition B.4 (Fractional Hypertree Width).
@@
Conjecture B.6 (OuMv (Henzinger et al., 2015)).
@@
Appendix C Missing Details in Section 5
In this section, we explain in detail our results on IVM in the insert-only setting, where no deletes are allowed. The main goal is to prove Theorem 3.1 (as well as Theorem 3.3). The results in this section also serve as building blocks for our results on the insert-delete setting in Section E.
C.1. Preliminaries: The Query Decomposition Lemma
We prove the query decomposition lemma (Lemma 26) and illustrate it by an example. This lemma is the core of the proof of Lemma 5.1, which in turn is a key ingredient in the proof of Theorem 3.1. Recall notation from Section B.
@@ C.1. Preliminaries: The Query Decomposition Lemma
Lemma C.1 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
@@
$\sum_{y\in\text{\sf Dom}(Y)}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|\sigma_{Y=y}R(\bm{X})|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
@@
$\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}\left(\sum_{y\in\text{\sf Dom}(Y)}|\sigma_{Y=y}R(\bm{X})|\right)^{\lambda_{R(\bm{X})}}$ $=$
$\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
@@
C.2. Proof of Lemma 5.1
Theorem 3.1 gives an upper bound on the amortized update time for IVM in the insert-only setting in terms of the fractional hypertree width of the query $Q$. Before we show how to achieve this bound in Sections C.3 and C.4, we prove in this section a weaker bound that is based on the AGM-bound of $Q$. In particular, we prove Lemma 5.1, which says the following. Given a query $Q$ (whose input relations are initially empty) and a stream of $N$ single-tuple inserts into $Q$, we can update $Q$ after every insert and report the new output tuples $\delta Q$ in a total time for all updates/reports together that does not exceed the AGM bound of $Q$. Note that here we are referring to the final AGM bound of $Q$ after all inserts have been performed.
@@
$\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(\tau)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$ $\leq\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$
$\leq\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}|^{\lambda_{R(\bm{X})}}=\theta$
@@
C.3. Upper bound for $\text{IVM}^{+}[Q]$
@@
For any query $Q$, both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(N^{\mathsf{w}(Q)-1})$ amortized update time, where $N$ is the number of single-tuple inserts.
@@
The case of $\text{IVM}^{+}[Q]$: Let $(\mathcal{T},\chi)$ be a tree decomposition of $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$. (Recall notation from Section B.) Fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. Given a node $t\in V(\mathcal{T})$, we will use $\text{\sf parent}(t),\text{\sf children}(t),\text{\sf ancestors}(t)$ and $\text{\sf decendants}(t)$ to refer to the parent, children, ancestors and descendants of $t$ in $\mathcal{T}$, respectively, under the fixed orientation of $\mathcal{T}$. For every node $t\in V(\mathcal{T})$, let $\bm{X}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\chi(t)$ be the variables of $t$ and $Q_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{\bm{X}_{t}}$ be the restriction of $Q$ to $\bm{X}_{t}$, given by Definition B.3. We call $Q_{t}$ a bag query. Let $\bm{X}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf vars}(Q)$. By definition, the query $Q$ is equivalent to the join of the outputs of all bag queries:
@@
C.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Theorem 3.1 for the $\text{IVM}^{+}_{\delta}[Q]$ problem. In particular, we explain how to modify the algorithm and proof from Section C.3, which was originally meant for the $\text{IVM}^{+}[Q]$ problem. Just like we did in Section C.3, here also we will construct an optimal tree decomposition of $Q$ that minimizes the fractional hypertree width, and we will maintain materialized relations for every bag of the tree decomposition. However, suppose that we have an insert of a tuple $\bm{t}$ into a relation $R$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this particular insert. We can no longer start our constant-delay enumeration from the root of the tree decomposition, as we did in Section C.3. Instead, we have to find the bag $B$ that contains $R$, select $\bm{t}$, and start our enumeration from $B$, as if $B$ were the root of the tree decomposition. In order to be able to enumerate the output of $Q$ starting from any bag $B$ of the tree decomposition, all bags need to be fully calibrated with each other. In particular, in Section C.3, we only calibrated bags in a bottom-up fashion where each bag was calibrated with its children bags. But here, we have to do full calibration consisting of two phases: One bottom-up phase where we calibrate each bag with its children, followed by a top-down phase where we calibrate each bag with its parent. These two calibrations are similar to the two phases of the Yannakakis algorithm (Yannakakis, 1981).
@@ C.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
For any query $Q$, both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(N^{\mathsf{w}(Q)-1})$ amortized update time, where $N$ is the number of single-tuple inserts.
@@
The case of $\text{IVM}^{+}_{\delta}[Q]$: We will use the same notation as in the proof of Theorem 3.1 for $\text{IVM}^{+}[Q]$ in Section C.3. In particular, here also we will fix an optimal tree decomposition $(\mathcal{T},\chi)$ of $Q$ and we will fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. In addition to $Q^{\prime}_{t}$ and $P^{\prime}_{t}$ given by (31) and (32), we will maintain two additional materialized relations $Q_{t}^{\prime\prime}$ and $P_{t}^{\prime\prime}$ for every node $t\in V(\mathcal{T})$. However, this time, we will define them inductively starting from the root $r$ all the way to the leaves of $\mathcal{T}$: (Recall that for the root $r$, we have $\bm{Y}_{r}=\emptyset$.)
@@
Now suppose we have an insert of a tuple $\bm{t}$ into a relation $R(\bm{Z})$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this insert. We find the node $t$ in $\mathcal{T}$ where $\bm{Z}\subseteq\chi(t)$. The existence of such a node $t$ is guaranteed by the definition of a tree decomposition (Definition B.2). Now, we re-orient the tree $\mathcal{T}$ so that $t$ becomes the root. We enumerate output tuples $\bm{x}_{t}$ of $Q^{\prime\prime}_{t}$ that join with $\bm{t}$. For each such tuple $\bm{x}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime\prime}_{c}$ for every child $c$ of $t$. (These are the children of $t$ under the new orientation of the tree $\mathcal{T}$). By (34), we guarantee that for each tuple $\bm{x}_{t}$ and each child $c$, there must be at least one tuple $x_{c}$ in $Q^{\prime\prime}_{c}$ that joins with $\bm{x}_{t}$. We continue this process inductively all the way to the leaves of $\mathcal{T}$. ∎
Example C.3 (for Theorem 3.1).
@@
C.5. Lower bound for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$
@@
For any query $Q$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ amortized update time, for any constant $\gamma>0$.
@@
Appendix D Missing Details in Section 6
In this section, we review some necessary background on intersection joins in the static setting. Let $Q$ be a query, $\overline{Q}$ be its time extension, and $\widehat{Q}$ be its multivariate extension. Suppose we want to evaluate $\overline{Q}$ over a given database instance $\overline{\mathcal{D}}$. This evaluation problem can be reduced (Khamis et al., 2022) to the evaluation of $\widehat{Q}$ over some database instance, $\widehat{\mathcal{D}}$, that satisfies $|\widehat{\mathcal{D}}|=\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$ and can be constructed in time $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$. Moreover, this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown (Khamis et al., 2022) by a backward reduction from each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail below. Recall that both reductions are in the context of static evaluation. However later on in Section E, we will rely on both reductions in order to prove upper and lower bounds on the IVM problem in the insert-delete setting.
D.1. A reduction from $\overline{Q}$ to $\widehat{Q}$
@@
Example D.2.
Suppose we want to solve the query $\overline{Q}_{\triangle}$ from Example 6.3. We can start by constructing a segment tree for the interval variable $[Z]$ and replacing each value $[z]$ of $[Z]$ by the components in its canonical partition $\text{\sf CP}_{N}([z])$. We can then use Lemma D.1 to look for intersections. In particular, let’s first try to find triples of segment-tree nodes $(z^{R},z^{S},z^{T})$ from $R,S,T$ respectively that appear on the same root-to-leaf path in the specific order $(z^{R},z^{S},z^{T})$. This can only happen if the bitstring $z^{R}$ is a prefix of $z^{S}$ which in turn is a prefix of $z^{T}$. To test this prefix pattern, we can break each $z^{S}$ into a concatenation two strings $z^{S}=z^{S}_{1}\circ z^{S}_{1}$ in all possible ways and similarly break $z^{T}$ into a concatenation of three string $z^{T}=z^{T}_{1}\circ z^{T}_{2}\circ z^{T}_{3}$ and then check whether $z^{R}=z^{S}_{1}=z^{T}_{1}$ and $z^{S}_{2}=z^{T}_{2}$. In particular, we construct the following input relations:
@@
$\widehat{T}_{123}$ =$=\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
@@
(37) $\widehat{Q}_{123}(Z_{1},Z_{2},Z_{3},A,B,C)=\widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
@@
Definition D.3 (Canonical partition of a relation $\text{\sf CP}_{N}^{(i)}(R)$).
@@
(39) $\text{\sf CP}_{N}^{(i)}([z],x_{1},\ldots,x_{m})\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},\ldots,z_{i},x_{1},\ldots,x_{m})\mid(z_{1}\circ\cdots\circ z_{i})\in\text{\sf CP}_{N}([z])\}$
@@
Definition D.4 (Canonical partition of a database instance $\text{\sf CP}_{N}(\overline{\mathcal{D}})$).
@@
Theorem D.5 (Implicit in (Khamis et al., 2022)).
@@
(41) $G_{k}(z_{1},\ldots,z_{k},x_{1},\ldots,x_{m})\stackrel{{\scriptstyle\text{def}}}{{=}}(z_{1}\circ\cdots\circ z_{k},x_{1},\ldots,x_{m})$
@@
D.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
In the previous Section D.1, we saw how to reduce the static evaluation of $\overline{Q}$ to $\widehat{Q}$. We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Khamis et al., 2022). Later on, we will use this backward reduction to prove lower bounds on the IVM problem in Section E.4.
Example D.6.